HUBBLE SPACE TELESCOPE DAILY REPORT #1621 PERIOD COVERED: 0000Z (UTC) 04/25/96 - 0000Z (UTC) 04/26/96 Daily Status Report as of 117/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 GHRS/1/2 6291 (Uranus: Lyman-Alpha and H2 Survey) The High Resolution Spectroscope (Sides One and Two) were used to = make ultraviolet observations of the planet Uranus, and its moon Titania. = This is to search for Lyman-Alpha and molecular hydrogen emission, as = well as other species generated by ion chemistry in the intense aurora = and ionosphere. A wavelength calibration image was also taken. The = observations were executed as scheduled, and no anomalies were noted. 1.2 Completed FGS 6262 (High Speed Astrometry - Barnard's Star) The Fine Guidance Sensors were used to make high speed astrometric = observations of Barnard's Star and seven reference stars. The = observations were executed as scheduled, and no anomalies were noted. 1.3 Completed WF/PC-2 5997 (The Cores of Elliptical Galaxies in the = Coma Cluster) The WF/PC-2 was used to make observations of the large elliptical = galaxy GAL-130009 which is located in the Coma Cluster. The purpose is = to study the luminosity profiles in the central nuclear region. The = observations were completed as scheduled, and no anomalies were noted. 1.4 Completed FOC/96 6057 (Ultra--High Resolution Studies of AGNs II: = 3C279 Proper Motion) The Faint Object Camera (f/96) was used to make ultra-high = resolution pictures of active galactic nuclei in radio galaxy 3C279 in = order to study the small scale morphology of the AGN. The data will be = searched for evidence of superluminal motion, which is known radio = observations, and for time variable phenomena. The observations were = executed as scheduled, and no anomalies were noted. 1.5 Completed WF/PC-2 5994 (HST Images of New Candidates For Distant = Proto-Ellipticals) The WF/PC-2 was used to make an observation of blue compact galaxy = SA57-17731 in order to measure its size, light profiles, and color. The = observation was executed unevenfully as scheduled. 1.6 Completed Two Sets of FOS/BL 6166 (FOS Cycle 5: Location of = Spectra and Wavelength Calibration) The Faint Object Spectrograph (Blue Side) was used to make = wavelength calibration images to measure the stability of the FOS Blue = Side wavelength scale and spectral locations. The observations were = completed as planned, and no anomalies were reported. 1.7 Completed Two Sets of WF/PC-2 6359 (Core Properties of the Bulges = of Spiral Galaxies) The WF/PC-2 was used to make snapshot survey observations of = spiral galaxy NGC-3900 in order to study its central bulges. The aim is = to investigate the core properties, correlation's with global galactic = properties, and the frequency of occurrence of multiple and/or = miscentered nuclei. The observations were completed as scheduled, and no = anomalies were noted. 1.8 Completed FOS/BL 6202 (Cycle 5 FOS Spectral Flat Field = Calibration) The Faint Object Spectrograph (Blue Side) was used to make = observations of the relatively featureless white dwarf star WD0501+527 in = order to make flat field calibration images. The observations were made = through the 1.0 aperture, with all usable detector/disperser = combinations. As described in 2.1 below, there was a problem with the = initial acquisition in this 10.5 hour proposal. HSTAR 5582 was also = written to document an FOS status buffer message at 116/191419Z = indicating that the HST was not considered stable, and that the take data = flag was down. Operations personnel saved the remainder of the proposal = by commanding the FOS door open at 116/2304Z. Otherwise, the remaining = observations were executed with no further problems. 1.9 Completed WF/P-2 6254 (GTO Pure Parallel Program - Cycle 5) The WF/PC-2 was use to make random long exposures of high galactic = fields taken in parallel with the prime science observations to obtain a = random sample of galaxies. The observations were executed as scheduled wi= th no anomalies. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 11 Successful : 10 As documented in HSTAR 5583, the acquisition at 116/190119Z = initially failed when the search radius was exceeded. The second attempt = also failed for the same reason. Subsequently, the return to RGA control = also failed, possibly due to conflict with the on-going FOS dwell scan. = The re-acquisition at 116/2032Z was successful. Scheduled Reacquisitions : 7 Successful : 7 2.2 FHST Updates: Scheduled : 18 Successful : 18 2.3 Operations Notes: There were no significant operational problems. 3.0 SIGNIFICANT FORTHCOMING EVENTS: The dry run for SMGT-22, Part 1 is being run today. /DMH